The average cholesterol content of a certain canned goods is 215 milligrams, and the standard deviation is 15 milligrams. Assume the variable is normally distributed:


a. If a canned goods is selected, what is the probability that the cholesterol content will be greater than 220 milligrams?


b. If a sample of 25 canned goods is selected, what is the probability that the mean of the samle will be larger than 220 milligrams?


1
Expert's answer
2021-03-30T07:37:37-0400

a. X ~ "N(215, 15^2)"

"P(X>220) = P(\frac{x-μ}{σ} > \frac{220-215}{15}) \\

= P(Z > 0.333) \\

= 1 -P(Z < 0.333) \\

= 1 -0.6304 \\

= 0.3696 \\

= 36.96 \; \%"

b. n = 25

"P(\bar{X}>220) = 1 -P(X<220) \\

= 1 -P(Z< \frac{x-μ}{\frac{σ}{\sqrt{n}}}) \\

= 1 – P(Z< \frac{220-215}{\frac{15}{\sqrt{20}}}) \\

= 1 -P(Z<1.49) \\

= 1 -0.9318 \\

= 0.068 \\

= 6.82 \; \%"